A train stops at two station P & Q which are 2 km apart .It accelerates uniformly from P at 1m/s*2 for 15sec & maintains a constant speed for a time before decelerating uniformly to rest at Q the decelration is 0.5ms*2, find time for which train is travelling at constant speed

Grade:11

1 Answers

Kruthic
6 years ago
When the train starts at P
Displacement in 15 seconds=ut+0.5at=0.5*1*15*15=225/2 metres.
Velocity attained=u+at=1*15=15m/s.
Let the train travel with 15m/s for t seconds.
Displacement covered by the train while travelling with constant velocity=vt=15t
While the train decelarates, v2-u2=2as
s=(v2-u2)/2a=-15*15/2(-0.5)=225metres
Total displacement=225/2+15t+225
2000=225/2+225+15t
15t=1662.5
Therefore , t=1662.5/15=110.83seconds.